A pool has a volume of 1000 ?3. At time ? = 0 the pool is empty. We then fill in water
the pool at a constant speed ? = 2.4?3 per minute. At the same time, water is leaking out
the pool at a rate that is at all times proportional to the volume of water.

The proportionality constant is ? = 3 ∙ 10−3??nute-1.

Let ? (?) be the volume of water in the pool t minutes after we started filling water in the pool.

a) Set up a differential equation that ? (?) must satisfy.
b) Solve the equation by the method with integrating factor.
c) Find when the pool is half full.
Answer #1

Assame that time ti amount of water in tanke = V Rate at which water enters Rin = a.4m m3/min Rate at which water leaves - Rodv 2 3x10 V- 2.4 du + at du [ di + at spat .3 S 3x10 at 3x10 3x10 t PCU) v= QCV) SP at Integrating factor.l.Fee Hege, entegr3x10 3x10 t v.e 800 e + C vCt) = 800+ cë 3x1õe when t=0, V=0 0 = 800+C > C=- 800 Solution VG+) = 800 C1- ē3xlot) = 800 C1- 7